# Project Page: https://lizhiqi49.github.io/MVControl/

### v2

@@ 5 Conclusion
In this work, we delve into the important yet under-explored field of controllable 3D generation. We present a novel network architecture, MVControl, for controllable text-to-multiview image generation. Our approach features a trainable control network that interacts with the base image diffusion model to enable controllable multi-view image generation. Once it is trained, Once trained, our network offers 3D diffusion guidance for controllable text-to-3D generation using a hybrid SDS gradient alongside another 2D diffusion model. We propose an efficient multi-stage 3D generation pipeline using both feed-forward and optimization-based methods. Our pioneering use of SuGaR—an explicit representation blending mesh and 3D Gaussians—outperforms previous Gaussian-based mesh generation approaches. Experimental results demonstrate our method’s ability to produce controllable, high-fidelity text-to-multiview images and text-to-3D assets. Furthermore, tests across various conditions show our method’s generalization capabilities. We believe our network has broader applications in 3D vision and graphics beyond controllable 3D generation via SDS optimization.
Appendix
Appendix 0.A Introduction
@@
Appendix 0.B Implementation Detail
0.B.1 Training Data
Multi-view Images Dataset. As described in our main paper, we filter the number of objects in Objaverse [14] to approximately 400k. For each retained sample, we first normalize its scene bounding box to a unit cube centered at the world origin. Subsequently, we sample a random camera setting by uniformly selecting the camera distance between 1.4 and 1.6, the angle of Field-of-View (FoV) between 40 and 60 degrees, the degree of elevation between 0 and 30 degrees, and the starting azimuth angle between 0 and 360 degrees. Under the random camera setting, multi-view images are rendered at a resolution of 256 under 4 canonical views at the same elevation starting from the sampled azimuth. During training, one of these views is chosen as the reference view corresponding to the condition image. We repeat this procedure three times for each object.
@@ 0.B.1 Training Data
0.B.2 Training Details of MVControl
While our base model, MVDream [50], is fine-tuned from Stable Diffusion v2.1 [45], we train our multi-view ControlNet models from publicly available 2D ControlNet checkpoints [^1] https://huggingface.co/thibaud adapted to Stable Diffusion v2.1 for consistency. The models are trained on an 8$\times$×A100 node, where we have 160 (40$\times$4) images on each GPU. With a gradient accumulation of 2 steps, we achieve a total batch size of 2560 images. We utilize a constant learning rate of $4\times 10^{-5}$ with 1000 steps of warm-up.
0.B.3 Implementation Details of 3D Generation
In our coarse Gaussian generation stage, the multi-view images are generated with our MVControl attached to MVDream using a 30-step DDIM sampler [51] with a guidance scale of 9 and a negative prompt "ugly, blurry, pixelated obscure, unnatural colors, poor lighting, dull, unclear, cropped, lowres, low quality, artifacts, duplicate". In the Gaussian optimization stage, the Gaussians are densified and pruned every 300 steps before step 1500. The 3D SDS ($\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$) is computed with a guidance scale of 50 using the CFG rescale trick [30], and $\nabla_{\theta}\mathcal{L}_{SDS}^{2D}$ is computed with a guidance scale of 20. The $\nabla_{\theta}\mathcal{L}_{VSD}$ in the SuGaR refinement stage is computed with a guidance scale of 7.5. These score distillation terms also incorporate the aforementioned negative prompt. Our implementation is based on the threestudio project [18]. All testing images for condition image extraction are downloaded from civitai.com.
Appendix 0.C Additional Qualitative Results
0.C.1 Diversity of MVControl
Figure 10: Diverse generated multi-view images of MVControl. Our MVControl can generate diverse multi-view images with same condition image and prompt.
Similar to 2D ControlNet [69], our MVControl can generate diverse multi-view images with the same condition image and prompt. Some of the results are shown in Fig. 10.
0.C.2 Textured Meshes
Figure 11: Additional qualitative results with Canny edge as input.
Figure 12: Additional qualitative results with normal map as input.
Figure 13: Additional qualitative results with depth map as input.
Figure 14: Additional qualitative results with user scribble as input.
We also provide additional generated textured mesh. The images under several views are shown in Figs. 11, 12, 13 and 14. Please refer to our project page for video and interactive mesh results.
Appendix 0.D Textual Prompts for 3D Comparison
@@
1. "RAW photo of A charming long brown coat dog, border collie, head of the dog, upper body, dark brown fur on the back,shelti,light brown fur on the chest,ultra detailed, brown eye"
2. "Wild bear in a sheepskin coat and boots, open-armed, dancing, boots, patterned cotton clothes, cinematic, best quality"
3. "Skull, masterpiece, a human skull made of broccoli"
4. "A cute penguin wearing smoking is riding skateboard, Adorable Character, extremely detailed"
5. "Masterpiece, batman, portrait, upper body, superhero, cape, mask"
6. "Ral-chrome, fox, with brown orange and white fur, seated, full body, adorable"
7. "Spiderman, mask, wearing black leather jacket, punk, absurdres, comic book"
8. "Marvel iron man, heavy armor suit, futuristic, very cool, slightly sideways, portrait, upper body"

### v3

@@ 5 Conclusion
In this work, we delve into the important yet under-explored field of controllable 3D generation. We present a novel network architecture, MVControl, for controllable text-to-multiview image generation. Our approach features a trainable control network that interacts with the base image diffusion model to enable controllable multi-view image generation. Once trained, our network offers 3D diffusion guidance for controllable text-to-3D generation using a hybrid SDS gradient alongside another 2D diffusion model. We propose an efficient multi-stage 3D generation pipeline using both feed-forward and optimization-based methods. Our pioneering use of SuGaR—an explicit representation blending mesh and 3D Gaussians—outperforms previous Gaussian-based mesh generation approaches. Experimental results demonstrate our method’s ability to produce controllable, high-fidelity text-to-multiview images and text-to-3D assets. Furthermore, tests across various conditions show our method’s generalization capabilities. We believe our network has broader applications in 3D vision and graphics beyond controllable 3D generation via SDS optimization.
Acknowledgement
This work was supported in part by NSFC under Grant 62202389, in part by a grant from the Westlake University-Muyuan Joint Research Institute, and in part by the Westlake Education Foundation.
Supplementary Material
Appendix A Introduction
@@
Appendix B Implementation Detail
B.1 Training Data
Multi-view Images Dataset. We employ the multi-view renderings from the publicly available large 3D dataset, Objaverse [14], to train our MVControl. Initially, we preprocess the dataset by removing all samples with a CLIP-score lower than 22, based on the labeling criteria from [52]. This filtering results in approximately 400k remaining samples. For each retained sample, we first normalize its scene bounding box to a unit cube centered at the world origin. Subsequently, we sample a random camera setting by uniformly selecting the camera distance between 1.4 and 1.6, the angle of Field-of-View (FoV) between 40 and 60 degrees, the degree of elevation between 0 and 30 degrees, and the starting azimuth angle between 0 and 360 degrees. Under the random camera setting, multi-view images are rendered at a resolution of 256$\times$256 under 4 canonical views at the same elevation starting from the sampled azimuth. We repeat this procedure three times for each object. During training, one of these views is chosen as the reference view corresponding to the condition image. Instead of utilizing the names and tags of the 3D assets, we employ the captions from [34] as text descriptions for our retained objects.
@@ B.1 Training Data
B.2 Training Details of MVControl
While our base model, MVDream [49], is fine-tuned from Stable Diffusion v2.1 [44], we train our multi-view ControlNet models from publicly available 2D ControlNet checkpointshttps://huggingface.co/thibaud adapted to Stable Diffusion v2.1 for consistency. The models are trained on an 8$\times$×A100 node, where we have 160 (40$\times$4) images on each GPU. With a gradient accumulation of 2 steps, we achieve a total batch size of 2560 images. The model undergoes 50000 steps of training under a constant learning rate of $4\times 10^{-5}$ with 1000 steps of warm-up. Similar to the approach in [68], we randomly drop the text prompt as empty with a 50% chance during training to facilitate classifier-free learning and enhance the model’s understanding of input condition images. Moreover, we also employ 2D-3D joint training following [49]. Specifically, we randomly sample images from the AES v2 subset of LAION [46] with a 30% probability during training to ensure the network retains its learned 2D image priors.
B.3 Implementation Details of 3D Generation
Multi-view Image Generation . In our coarse Gaussian generation stage, the multi-view images are generated with our MVControl attached to MVDream using a 30-step DDIM sampler [50] with a guidance scale of 9 and a negative prompt ”ugly, blurry, pixelated obscure, unnatural colors, poor lighting, dull, unclear, cropped, lowres, low quality, artifacts, duplicate”.
Gaussian Optimization Stage . This stage comprises a total of 3000 steps. During the initial 1500 steps, we perform simple 3D Gaussian optimization with split and prune every 300 steps. After step 1500, we cease densification and prune, and instead introduce SuGaR [17] regularization terms to refine the scene. In the end of the stage, we prune all Gaussians with opacity below $\bar{\sigma}=0.5$. The 3D SDS ($\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$) is computed with a guidance scale of 50 using the CFG rescale trick [29], and $\nabla_{\theta}\mathcal{L}_{SDS}^{2D}$ is computed with a guidance scale of 20. We use $\lambda_{2D}=0.1$ and $\lambda_{3D}=0.01$ for 2D and 3D diffusion guidance, with resolutions of 512$\times$512 and 256$\times$256 respectively.
SuGaR Refinement Stage . This stage has totally 5000 steps of optimiztion. The $\nabla_{\theta}\mathcal{L}_{VSD}$ in the SuGaR refinement stage is computed with a guidance scale of 7.5. The training is under resolution 512$\times$512 for rendering.
All score distillation terms also incorporate the aforementioned negative prompt. Our implementation is based on the threestudio project [18]. All testing images for condition image extraction are downloaded from civitai.com.
Appendix C Additional Qualitative Results
C.1 Diversity of MVControl
Similar to 2D ControlNet [68], our MVControl can generate diverse multi-view images with the same condition image and prompt. Please refer to our project page for some of the results.
C.2 Textured Meshes
We also provide additional generated textured mesh. Please refer to our project page for video and interactive mesh results.
Appendix D Textual Prompts for 3D Comparison
@@
1. ”RAW photo of A charming long brown coat dog, border collie, head of the dog, upper body, dark brown fur on the back,shelti,light brown fur on the chest,ultra detailed, brown eye”
2. ”Wild bear in a sheepskin coat and boots, open-armed, dancing, boots, patterned cotton clothes, cinematic, best quality”
3. ”Skull, masterpiece, a human skull made of broccoli”
4. ”A cute penguin wearing smoking is riding skateboard, Adorable Character, extremely detailed”
5. ”Masterpiece, batman, portrait, upper body, superhero, cape, mask”
6. ”Ral-chrome, fox, with brown orange and white fur, seated, full body, adorable”
7. ”Spiderman, mask, wearing black leather jacket, punk, absurdres, comic book”
8. ”Marvel iron man, heavy armor suit, futuristic, very cool, slightly sideways, portrait, upper body”